OpenStudy (anonymous):

Guys help! Determine the slope and the y-intercept of the line whose equation is given. 2y + 5x −8 = 0

OpenStudy (anonymous):

use the formula y = mx+b

OpenStudy (anonymous):

wherein m is your slope

OpenStudy (anonymous):

y = (-5/2) x +8

OpenStudy (ineedhelplz):

\[\frac{ Rise }{ Run }\] Is a tip that always used to help me :P

OpenStudy (anonymous):

oops it should be y = (-5/2) x +8/2

OpenStudy (anonymous):

y = -5/2x +4

OpenStudy (anonymous):

b is your y-intercept

OpenStudy (anonymous):

so your slope is -5/2 and your y-intercept is?

OpenStudy (xmissalycatx):

You want to get this into y = mx + b form to determine this! So what you need to do is: - add 8 to both sides so you have 2y + 5x = 8 -now you need to subtract 5x from both sides so you have 2y = -5x + 8 -next you need to divide EVERYTHING by 2! So now you'll get y = -5x over 2 + 4 m (or -5x/2) is your slope and your b is 4! Hope this helped! Have a lovely day! <3

OpenStudy (koikkara):

@Zaada Does it Help You.?

OpenStudy (anonymous):

@xMissAlyCatx Thank you so much! You helped me a looot :)

OpenStudy (anonymous):

@Koikkara Thaaaanks! It did help :)

OpenStudy (anonymous):

@martaamador62 Thank you!! I got it :)

OpenStudy (anonymous):

np :)